Find the exact values of sin(θ/2), cos(θ/2), and tan(θ/2) for the given conditions.
sec θ = −
4
3
; 180° < θ < 270°

Answer by stanbon(75887)

Find the exact values of sin(θ/2), cos(θ/2), and tan(θ/2) for the given conditions.
sec θ = −4/3 ;
Note:: cos(t) = -3/4
------------------------
x = -3, r = 4, y = -sqrt[4^2-3^2] = -sqrt(7)
-----
sin(t/2) = -sqrt[(1+3/4)/2] = -sqrt[7/8]
cos(t/2) = -sqrt[(1-(3/4)/2] = -sqrt[1/8]
tan(t/2) = -sqrt(7/8)/-sqrt(1/8) = sqrt[7]
180° < θ < 270°
--------
Cheers,
Stan H.
------------